a hospital nursery has 20 newborn babies, of which 8 are male. Two of the babies are taken home by their parents. What is the probability that at least one of the two is male? Note: Since you have no further information, the answer is the same as though the two babies were selected at random.
a.127/190 b. 48/95 c. 5/8 d.94/95 e.62/95
can you explain how to get the answer please?

Answer by sudhanshu_kmr(1152)


8 are male, 12 are female.
probability that both are female = 12/20 * 11/19 = 33/95
probability that at least one is male = 1 - 33/95 = 62/95

e. 62/95